A contractor mexes concrete from bags of pre-max for small jobs. How many bags with 10% cement should he mix with 8 bags of 12.5% cement to produce a mix containing 11% cement?
Answer by checkley75(3666)

.10x+.125*8=.11(x+8)
.10x+1=.11x+.88
.10x-.11x=.88-1
-.01x=-.12
x=-.12/-.01
x=12 bags of 10% cement is need.
proof
.10*12+.125*8=.11(12+8)
1.2+1=.11(20)
2.2=2.2